OpenStudy (anonymous):

what is the square root of x^2/5

OpenStudy (anonymous):

Think of x^(2/5) as [x^(1/5)]^2 Then you have the + and - square root of that: +- x^(1/5)

OpenStudy (zehanz):

So you have:\[\sqrt{\frac{ x^2 }{ 5 }}\]You can split up:\[\frac{ \sqrt{x^2} }{ \sqrt{5} }\]

OpenStudy (anonymous):

Are you looking for the square root of: x^(2/5) or (x^2)/5 ?

OpenStudy (anonymous):

the 2/5 is the exponent tho?

OpenStudy (anonymous):

its the first one :)

OpenStudy (anonymous):

If that is the case, then I stick with my first post.

OpenStudy (anonymous):

All good now?

OpenStudy (zehanz):

Well, apparently, you have x^(2/5). You might consider writing it that way yourself to avoid misunderstanding... ;) \[x^{\frac{ 2 }{ 5 }}=\sqrt[5]{x^2}=(\sqrt[5]{x})^2\]Take your pick!

OpenStudy (zehanz):

if you need the square root of x^(2/5) there is more to do: 1. with exponents only:\[(x^{\frac{ 2 }{ 5 }})^{\frac{ 1 }{ 2 }}=x^{\frac{ 1 }{ 5 }}=\sqrt[5]{x}\] 2. with the result of my last calculation:\[\sqrt{(\sqrt[5]{x})^2 }=\sqrt[5]{x}\] So the answer is the 5th root of x.